# Enhancing Text Editing for Grammatical Error Correction: Arabic as a Case Study

### v1

@@ Abstract
Bashar Alhafni  and Nizar Habash Computational Approaches to Modeling Language Lab New York University Abu Dhabi {alhafni,nizar.habash}@nyu.edu
@@
Bashar Alhafni  and Nizar Habash Computational Approaches to Modeling Language Lab New York University Abu Dhabi {alhafni,nizar.habash}@nyu.edu
@@ 1 Introduction
Grammatical Error Correction (GEC) is a well-studied problem, particularly in English, with numerous datasets and shared tasks Ng et al. (2013, 2014); Bryant et al. (2019). GEC has applications in both writing assistance for native speakers (L1) and language learning for second-language (L2) learners. While neural machine translation (NMT) approaches have long dominated GEC and continue to achieve strong results when trained on large amounts of data Stahlberg and Kumar (2024); Bryant et al. (2023), they are not inherently the most efficient. Unlike MT, where input and output sequences differ significantly, GEC typically involves minimal changes, with most input tokens copied to the output. Employing full-sequence autoregressive models in such cases can be computationally wasteful Stahlberg and Kumar (2020). A highly efficient and competitive alternative to sequence-to-sequence (Seq2Seq) models is text editing, which frames GEC as a sequence tagging problem. Instead of generating text autoregressively, text editing models assign edit labels to input tokens, leading to a more efficient and interpretable corrections. However, most commonly used text editing approaches require effort to design language-specific edit tag sets Awasthi et al. (2019); Omelianchuk et al. (2020); Mesham et al. (2023). This limits their adaptability for morphologically rich languages like Arabic Kwon et al. (2023), where the number of possible edits can be vast. Inspired by recent advancements in text editing Awasthi et al. (2019); Malmi et al. (2019); Omelianchuk et al. (2020); Straka et al. (2021); Mesham et al. (2023), we introduce a novel text editing approach that eliminates the need for language-specific edits. Instead, our method derives edit tags directly from data, making it more adaptable and scalable across different linguistic settings. We demonstrate the effectiveness of our approach on Arabic GEC. Our contributions are as follows:
@@ 2.2 Arabic Grammatical Error Correction
The first major efforts on MSA GEC were initiated by the Qatar Arabic Language Bank (QALB) project Zaghouani et al. (2014, 2015), which organized the QALB-2014 Mohit et al. (2014) and QALB-2015 Rozovskaya et al. (2015) shared tasks. More recently, Habash and Palfreyman (2022) introduced the ZAEBUC corpus, a dataset of essays written by native Arabic-speaking university students. Approaches to MSA GEC have included feature-based classifiers Rozovskaya et al. (2014); Farra et al. (2014); Bougares and Bouamor (2015); Nawar (2015) and NMT-based systems Watson et al. (2018); Solyman et al. (2021, 2022, 2023). LLMs have also been evaluated for MSA GEC Kwon et al. (2023), but attempts to adapt text editing models have been largely ineffective. The current SOTA was established by Alhafni et al. (2023), who incorporated contextualized morphological preprocessing and grammatical error detection (GED) features into Seq2Seq models, achieving SOTA results on the QALB-2014, QALB-2015, and ZAEBUC datasets.
@@ 3.2 Edit Representation
3.3 Edits Coverage
Table 1 presents edit statistics for QALB-2014, illustrating the impact of our strategies to reduce the edit vocabulary size $|E|$ on edit coverage and upper-bound (oracle) performance on the development (Dev) set. Edit coverage measures the proportion of training edits found in the Dev set, while oracle performance is evaluated using the MaxMatch (M2) scorer Dahlmeier and Ng (2012) F0.54.2). We use AraBERTv02 Antoun et al. (2020) for subword tokenization, as it yielded the best results among our tested models (more details in §5). Switching from word-level to subword-level edits reduces unique training edits by 44% (16,221 to 9,060) and lowers the Dev set OOV rate from 1% to 0.4%, yielding a 0.3-point F0.5 gain. Edit compression further reduces unique edits while preserving OOVand oracle performance. Segregating punctuation (Pnx) from non-punctuation (NoPnx) edits reduces combined training edits (4,799+160 from 6,170). However, NoPnx results are not directly comparable since punctuations are explicitly removed before the evaluation. Pnx F0.5 scores are higher as they are evaluated on a Dev set with non-punctuation errors already corrected, making the test easier. To assess the impact of pruning, we apply frequency thresholds of 10, 20, and 30 to remove low-frequency edits. As expected, pruning reduces the number of unique training edits and increases the OOVin the Dev set, yet F0.5 remains largely unaffected. This suggests that the majority of the 6,170 compressed subword edits occur infrequently and contribute little to the model’s upper-bound performance. A similar trend is observed for both Pnx and NoPnx edits, reinforcing the idea that many low-frequency edits can be pruned without degrading oracle performance. We present the same analysis on all datasets in Appendix B Table 8.
@@ 4.1 Data
We report on three publicly available MSA GEC datasets. The first is the QALB-2014 shared task dataset Mohit et al. (2014), followed by the native (L1) test set from the QALB-2015 shared task Rozovskaya et al. (2015). The third dataset is ZAEBUC Habash and Palfreyman (2022). Both QALB-2014 and the L1 test set of QALB-2015 contain native user comments from the Aljazeera news website, whereas ZAEBUC consists of essays written by native Arabic-speaking university students. We use the publicly available splits for QALB-2014 and QALB-2015, while for ZAEBUC, we use the splits created by Alhafni et al. (2023).
@@
We use the MADAR CODA corpus Eryani et al. (2020), a collection of 10,000 sentences from five Arabic city dialects (Beirut, Cairo, Doha, Rabat, and Tunis) represented in the CODA standard in parallel with their original raw form. The sentences come from the Multi-Arabic Dialect Applications and Resources (MADAR) Project Bouamor et al. (2018) and are in parallel across the cities (2,000 sentences from each city). We use the publicly available splits created by Alhafni et al. (2024). Table 2 summarizes the dataset statistics.
@@ 4.2 Evaluation
MADAR CODA Train 7K 25K 24% Comments
Dev 1.5K 5K 25% Comments
Test 1.5K 5K 26% Comments
Table 2: Corpus statistics of MSA and DA GEC datasets.
@@
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 83.7 68.8 75.6 80.3 86.7 73.9 79.8 83.8
@@
+ GPT-4o 89.1 61.6 72.8 81.8 93.3 68.3 78.9 86.9
Table 3: MSA GEC results on the Dev sets of QALB-2014 and ZAEBUC. A’2023 is Alhafni et al. (2023). Best non-ensemble results are underlined. The best overall results are in bold.
@@
+ GPT-4o 93.8 72.5 81.8 88.6
Table 4: DA GEC results on MADAR CODA Dev set. A’2024 is Alhafni et al. (2024). Best non-ensemble results are underlined, best overall results are in bold.
@@ 4.3 Models
We evaluate four LLMs: two commercial models and two open-source, Arabic-centric models. The commercial models include OpenAI’s GPT-3.5-turbo and GPT-4o OpenAI et al. (2024), while the Arabic-centric models are Jais-13B-Chat Sengupta et al. (2023) and the recently introduced Fanar LLM Team et al. (2025). We prompt GPT-3.5-turbo, GPT-4o, and Fanar through the OpenAI API, while Jais-13B-Chat is prompted using Hugging Face’s Transformers Wolf et al. (2020). Our experiments use both English and Arabic prompts, employing 0-shot and 5-shot prompting strategies. We design the prompts to elicit minimal edit-style corrections, ensuring that the LLMs’ outputs remain as close as possible to the original input in phrasing and lexical choices. We present our prompts in Figures 2 and 3 in Appendix E.
@@
To investigate the impact of the edit representation design on performance (§3.2), we build several edit taggers with different configurations. For word-level tagging, we use the representation of the first subword of each word and pass it through the subsequent layers. For subword-level tagging, we use the representation of each subword individually. Several Arabic pretrained transformer encoders based on BERT Devlin et al. (2019) have been developed Antoun et al. (2020); Abdul-Mageed et al. (2021); Inoue et al. (2021); Ghaddar et al. (2022). We select the three best-performing Arabic BERT models, as identified by Inoue et al. (2021) across various sentence and token classification tasks: AraBERTv02 Antoun et al. (2020), ARBERTv2 Abdul-Mageed et al. (2021), and CAMeLBERT-MSA Inoue et al. (2021). For QALB-2014, our edit taggers are trained exclusively on QALB-2014, following the shared task restrictions. For QALB-2015 (L1), we train only on QALB-2014 for consistency. For ZAEBUC, we train on both QALB-2014 and ZAEBUC, upsampling ZAEBUC tenfold to address its smaller size and domain shift. For DA GEC, we train on the MADAR CODA training split. The hyperparameters we used are detailed in Appendix A.
@@
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 84.5 67.7 75.2 80.5 82.2 73.6 77.7 80.3 85.7 74.1 79.5 83.1
@@
+ GPT-4o 89.7 60.2 72.0 81.7 88.3 66.7 76.0 82.9 93.4 68.9 79.3 87.2
Table 5: MSA GEC results on the Test sets of QALB-2014, QALB-2015 (L1), and ZAEBUC. A’2023 is Alhafni et al. (2023). Best non-ensemble results are underlined. The best overall results are in bold.
@@
3-Ensemble 92.2 77.7 84.3 88.9
+ GPT-4o 94.0 72.9 82.1 88.8
Table 6: DA GEC results on the Test set of MADAR CODA. A’2024 is Alhafni et al. (2024). Best non-ensemble results are underlined, best overall results are in bold.
@@ 5 Results
Tables 3 and 4 show the Dev results for MSA and DA GEC, respectively. For MSA GEC, we compare our results with those of Alhafni et al. (2023), including their best Seq2Seq models and their Seq2Seq++ setup, which incorporates morphological preprocessing and GED information. For DA GEC, we compare with Alhafni et al. (2024), including their best Seq2Seq model and Seq2Seq++ setup, which integrates dialect identification.
@@
We present LLMs results on MSA and DA GEC using their best setups, optimized for average F0.5 across all datasets based on prompt language and strategy (0-shot vs. 5-shot). Full results are in Table 10 (Appendix D). For QALB-2014, GPT-4o and Fanar outperform GPT-3.5 and Jais-13B-Chat, with GPT-4o achieving the best performance, though none surpass Alhafni et al. (2023). On ZAEBUC, GPT-4o leads, achieving the highest recall (76.8) and F1 (81.3). For DA GEC, GPT-4o remains the top LLM, but overall LLM performance is notably lower than for MSA.
@@
The full edit tagging results on the Dev sets, exploring edit design choices with CAMeLBERT-MSA, AraBERTv02, and ARBERTv2, are in Table 9 (Appendix C). AraBERTv02 consistently performs best. Subword-level edits, compression, and pruning improve performance, with optimal pruning thresholds of 10 for QALB-2014 and MADAR CODA, and 30 for ZAEBUC. The optimal setup for each dataset (subword, compression, pruning) is presented in Tables 3 and 4. We henceforth refer to this system as sweet (Subword Edit Error Tagger). sweet achieves an F0.5 of 78.8 on QALB-2014 and 86.0 on MADAR CODA, outperforming the Seq2Seq baseline on QALB-2014 and setting a new SOTA on MADAR CODA. On ZAEBUC, it scores 82.7 F0.5, trailing behind the Seq2Seq baseline. Consistent with previous work on text editing Omelianchuk et al. (2020); Straka et al. (2021), we find that iterative correction improves MSA GEC up to two iterations ($\textsc{sweet}^{2}$), achieving 79.3 on QALB-2014 and 83.0 F0.5 on ZAEBUC, with the highest overall recall on QALB-2014 (70.4). However, iterative correction degrades DA GEC performance. Separating non-punctuation edits ($\textsc{sweet}_{\text{NoPnx}}$) from punctuation edits ($\textsc{sweet}_{\text{Pnx}}$) improves MSA GEC performance. The best setup applies these systems in sequence: two iterations of non-punctuation correction followed by one iteration of punctuation correction ($\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$). This setup achieves the highest F0.5 score among text-editing models, setting a new SOTA on QALB-2014 with 80.3 (driven by a precision improvement of 83.7). Its performance on ZAEBUC leads other edit tagging techniques but trails behind GPT-4o. For our ensemble models (3-Ensemble), we combine the top three non-LLM models per dataset. For QALB-2014 and ZAEBUC, we use Seq2Seq++, $\textsc{sweet}^{2}$, and the cascaded setup $\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$. For MADAR CODA, we combine Seq2Seq++, sweet, and the second-best sweet model using CAMeLBERT-MSA (see Table 9 in Appendix C). The 3-Ensembles outperform single models, achieving SOTA results across all datasets, primarily through increased precision at the cost of recall. Adding GPT-4o further boosts performance (+GPT-4o), reaching an F0.5 of 81.8 on QALB-2014, 86.9 on ZAEBUC, and 88.6 on MADAR CODA.
@@
Tables 5 and 6 present the Test results for MSA and DA GEC, using the best setups identified from the Dev sets. On QALB-2014, the cascaded setup $\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ sets a new SOTA with 80.5 F0.5, outperforming the Seq2Seq and Seq2Seq++ models from Alhafni et al. (2023). On QALB-2015, this setup matches the Seq2Seq++ baseline with an F0.5 of 80.3. Similarly, on ZAEBUC, it achieves 83.1, on par with Seq2Seq++. Our ensemble models further enhance performance across all datasets, with the best results reaching 81.7 on QALB-2014, 82.9 on QALB-2015, 87.2 on ZAEBUC, and 88.9 on MADAR CODA in terms of F0.5. It is worth noting that adding GPT-4o to the ensemble boosts MSA GEC performance but does not improve DA GEC.
5.1 Runtime Performance

### v2

@@ Abstract
[…]
@@
[…]
@@ 1 Introduction
Grammatical Error Correction (GEC) is a well-studied problem, particularly in English, with numerous datasets and shared tasks Ng et al. (2013, 2014); Bryant et al. (2019). GEC has applications in both writing assistance for native speakers (L1) and language learning for second-language (L2) learners. While neural machine translation (NMT) approaches have long dominated GEC and continue to achieve strong results when trained on large amounts of data Stahlberg and Kumar (2024); Bryant et al. (2023), they are not inherently the most efficient. Unlike MT, where input and output sequences differ significantly, GEC typically involves minimal changes, with most input tokens copied to the output. Employing full-sequence autoregressive models in such cases can be computationally wasteful Stahlberg and Kumar (2020). A highly efficient and competitive alternative to sequence-to-sequence (Seq2Seq) models is text editing, which frames GEC as a sequence tagging problem. Instead of generating text autoregressively, text editing models assign edit labels to input tokens, leading to a more efficient and interpretable corrections. However, most popular text editing approaches require effort to design language-specific edit tag sets Awasthi et al. (2019); Omelianchuk et al. (2020); Mesham et al. (2023). This limits their adaptability for morphologically rich languages like Arabic Kwon et al. (2023), where the space of possible edits is large. Inspired by recent advancements in text editing Awasthi et al. (2019); Malmi et al. (2019); Omelianchuk et al. (2020); Straka et al. (2021); Mesham et al. (2023), we introduce a novel text editing approach that eliminates the need for language-specific edits. Instead, our method derives edit tags directly from data, making it more adaptable and scalable across different linguistic settings. We demonstrate the effectiveness of our approach on Arabic GEC. Our contributions are as follows:
@@ 2.2 Arabic Grammatical Error Correction
The first major efforts on MSA GEC were initiated by the Qatar Arabic Language Bank (QALB) project Zaghouani et al. (2014, 2015), which organized the QALB-2014 Mohit et al. (2014) and QALB-2015 Rozovskaya et al. (2015) shared tasks. More recently, Habash and Palfreyman (2022) introduced the ZAEBUC corpus, a dataset of essays written by native Arabic-speaking university students. Approaches to MSA GEC have included feature-based classifiers Rozovskaya et al. (2014); Farra et al. (2014); Bougares and Bouamor (2015); Nawar (2015) and NMT-based systems Watson et al. (2018); Solyman et al. (2021, 2022, 2023). LLMs have also been evaluated for MSA GEC Kwon et al. (2023); Alhafni et al. (2023); Magdy et al. (2024), but attempts to adapt text editing models have been largely ineffective. The current SOTA was established by Alhafni et al. (2023), who incorporated contextualized morphological preprocessing and grammatical error detection (GED) features into Seq2Seq models, achieving SOTA results on the QALB-2014, QALB-2015, and ZAEBUC datasets.
@@ 3.2 Edit Representation
3.3 Edit Coverage
Table 1 presents edit statistics for QALB-2014, illustrating the impact of our strategies to reduce the edit vocabulary size $|E|$ on edit coverage and upper-bound (oracle) performance on the development (Dev) set. Edit coverage measures the proportion of training edits found in the Dev set, while oracle performance is evaluated using the MaxMatch (M2) scorer Dahlmeier and Ng (2012) F0.54.2). We use AraBERTv02 Antoun et al. (2020) for subword tokenization, as it yielded the best results among our tested models (more details in §5). Switching from word-level to subword-level edits reduces unique training edits by 44% (16,221 to 9,060) and lowers the Dev set OOV rate from 1% to 0.4%, yielding a 0.3-point F0.5 gain. Edit compression further reduces unique edits while preserving OOVand oracle performance. Segregating punctuation (Pnx) from non-punctuation (NoPnx) edits reduces combined training edits (4,799+160 from 6,170). However, NoPnx results are not directly comparable, as punctuation is explicitly removed before the evaluation. Pnx F0.5 scores are higher as they are evaluated on a Dev set with non-punctuation errors already corrected, making the test easier. To assess the impact of pruning, we apply frequency thresholds of 10, 20, and 30 to remove low-frequency edits. As expected, pruning reduces the number of unique training edits and increases the OOVin the Dev set, yet F0.5 remains largely unaffected. This suggests that the majority of the 6,170 compressed subword edits occur infrequently and contribute little to the model’s upper-bound performance. A similar trend is observed for both Pnx and NoPnx edits, reinforcing the idea that many low-frequency edits can be pruned without degrading oracle performance. We present the same analysis on all datasets in Appendix B Table 9.
@@ 4.1 Data
We report on three publicly available MSA GEC datasets. The first is the QALB-2014 shared task dataset Mohit et al. (2014), followed by the native (L1) test set from the QALB-2015 shared task Rozovskaya et al. (2015). The third dataset is ZAEBUC Habash and Palfreyman (2022). QALB-2014 and the L1 test set of QALB-2015 contain comments by native speakers from the Aljazeera news website, whereas ZAEBUC consists of essays written by native university students. We use the publicly available splits for QALB-2014 and QALB-2015, while for ZAEBUC, we use the splits created by Alhafni et al. (2023).
@@
We use the MADAR CODA corpus Eryani et al. (2020), a set of 10,000 sentences from five Arabic city dialects (Beirut, Cairo, Doha, Rabat, and Tunis) written in the CODA standard in parallel with their original raw form. The sentences come from the Multi-Arabic Dialect Applications and Resources (MADAR) Project Bouamor et al. (2018) and are in parallel across the cities (2,000 sentences per city). We use the publicly available splits created by Alhafni et al. (2024). Table 2 summarizes the dataset statistics.
@@ 4.2 Evaluation
MADAR CODA Train 7K 40K 22% Comments
Dev 1.5K 9K 20% Comments
Test 1.5K 9K 21% Comments
Table 2: Corpus statistics of MSA (QALB, ZAEBUC) and DA (MADAR CODA) GEC datasets.
@@
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 83.7 68.8 75.6 80.3$\dagger$ 86.7 73.9 79.8 83.8
@@
4-Ensemble 89.1 61.6 72.8 81.8$\ddagger$ 93.3 68.3 78.9 86.9$\ddagger$
Table 3: MSA GEC results on the Dev sets of QALB-2014 and ZAEBUC. A’2023 is Alhafni et al. (2023). Best non-ensemble results are underlined; best overall results are in bold. $\dagger$ denotes statistical significance over the best baseline; $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
@@
4-Ensemble 93.8 72.5 81.8 88.6$\ddagger$
Table 4: DA GEC results on the MADAR CODA Dev set. A’2024 is Alhafni et al. (2024). Best non-ensemble results are underlined; best overall results are in bold. $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
@@ 4.3 Models
We evaluate four LLMs: two commercial models and two open-source, Arabic-centric models. The commercial models include OpenAI’s GPT-3.5-turbo and GPT-4o OpenAI et al. (2024), while the Arabic-centric models are Jais-13B-Chat Sengupta et al. (2023) and the recently introduced Fanar LLM Team et al. (2025). We prompt GPT-3.5-turbo, GPT-4o, and Fanar through the OpenAI API, while Jais-13B-Chat is prompted using Hugging Face’s Transformers Wolf et al. (2020). Our experiments use both English and Arabic prompts, employing 0-shot and 5-shot prompting strategies. We design the prompts to elicit minimal edit-style corrections, ensuring that the LLMs’ outputs remain as close as possible to the original input in phrasing and lexical choices. We present our prompts in Figures 2 and 3 in Appendix H.
@@
To investigate the impact of edit representation design on performance (§3.2), we build edit taggers with different configurations. For word-level tagging, we use the representation of the first subword of each word and pass it through the subsequent layers. For subword-level tagging, we use the representation of each subword individually. Several Arabic pretrained transformer encoders based on BERT Devlin et al. (2019) have been developed Antoun et al. (2020); Abdul-Mageed et al. (2021); Inoue et al. (2021); Ghaddar et al. (2022). We select the three best-performing Arabic BERT models, as identified by Inoue et al. (2021) across various sentence and token classification tasks: AraBERTv02 Antoun et al. (2020), ARBERTv2 Abdul-Mageed et al. (2021), and CAMeLBERT-MSA Inoue et al. (2021). For QALB-2014, our edit taggers are trained exclusively on QALB-2014, following the shared task restrictions. For QALB-2015 (L1), we train only on QALB-2014 for consistency. For ZAEBUC, we train on both QALB-2014 and ZAEBUC, upsampling ZAEBUC tenfold to address its smaller size and domain shift. For DA GEC, we train on the MADAR CODA training split. The hyperparameters we used are detailed in Appendix A.
@@
$\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ 84.5 67.7 75.2 80.5$\dagger$ 82.2 73.6 77.7 80.3 85.7 74.1 79.5 83.1
@@
4-Ensemble 89.7 60.2 72.0 81.7$\ddagger$ 88.3 66.7 76.0 82.9$\ddagger$ 93.4 68.9 79.3 87.2$\ddagger$
Table 5: MSA GEC results on the Test sets of QALB-2014, QALB-2015 (L1), and ZAEBUC. A’2023 is Alhafni et al. (2023). Best non-ensemble results are underlined; best overall results are in bold. $\dagger$ denotes statistical significance over the best baseline; $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
@@
3-Ensemble 92.2 77.7 84.3 88.9$\ddagger$
4-Ensemble 94.0 72.9 82.1 88.8
Table 6: DA GEC results on the Test set of MADAR CODA. A’2024 is Alhafni et al. (2024). Best non-ensemble results are underlined, best overall results are in bold. $\ddagger$ denotes statistical significance over both the best baseline and the best non-ensemble model.
@@ 5 Results
Tables 3 and 4 show the Dev results for MSA and DA GEC, respectively. For each dataset, we compare our models with the best-performing Seq2Seq and Seq2Seq++ baselines reported by Alhafni et al. (2023) and Alhafni et al. (2024). The Seq2Seq++ setups incorporate additional signals, such as morphological preprocessing and GED information for MSA GEC, or dialect identification for DA GEC. Full results for all Seq2Seq-based baseline variants across datasets are provided in Appendices E and F.
@@
We present LLMs results on MSA and DA GEC using their best setups, optimized for average F0.5 across all datasets based on prompt language and strategy (0-shot vs. 5-shot). Full results are in Table 11 (Appendix D). For QALB-2014, GPT-4o and Fanar outperform GPT-3.5 and Jais-13B-Chat, with GPT-4o achieving the best performance, though none surpass Alhafni et al. (2023). On ZAEBUC, GPT-4o leads, achieving the highest recall (76.8) and F1 (81.3). For DA GEC, GPT-4o is the top LLM, but overall LLM performance is notably lower than for MSA.
QALB-2014 ZAEBUC MADAR CODA
Baseline sweet Ensemble Baseline sweet Ensemble Baseline sweet Ensemble
Delete 41.1 44.3 45.2 51.9 63.6 62.5 0.0 0.0 0.0
Merge-B 94.0 93.7 93.8 96.7 96.9 96.6 94.4 86.6 92.7
Merge-I 93.8 93.5 93.6 96.7 96.9 96.6 93.6 84.8 91.6
M 33.9 33.6 28.6 48.6 50.0 41.7 82.5 78.0 82.4
M+O 58.0 61.0 60.6 55.6 100.0 0.0 0.0 0.0 0.0
O 94.3 94.5 94.4 94.4 94.4 94.1 92.1 90.2 91.4
O+X 78.1 81.5 83.3 0.0 0.0 0.0 0.0 0.0 0.0
P 75.0 75.6 76.8 62.8 71.9 70.4 65.8 35.7 55.6
S 46.4 57.1 57.4 40.4 47.6 46.9 83.1 82.3 83.0
X 61.2 61.4 62.4 72.9 74.1 74.1 73.8 76.9 79.5
Split 87.1 83.8 87.6 88.2 90.0 95.2 85.9 83.3 86.8
UNK 59.2 55.0 56.0 63.1 44.6 47.6 93.0 94.6 94.1
C 97.1 96.4 94.7 96.1 96.0 93.9 97.0 96.0 94.7
Macro Avg. 70.7 71.6 71.9 66.7 71.2 63.1 78.3 73.5 77.4
Table 7: Error type performance on the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA for the best Seq2Seq++ baseline, the best sweet model ($\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ for QALB-2014 and ZAEBUC; sweet for MADAR CODA), and the best ensemble (4-Ensemble). Results are reported in terms of F0.5. Best non-ensemble results are underlined; best overall results are in bold. UNK refers to unknown error types; C refers to correct words.
@@
Table 10 (Appendix C) presents the full edit tagging results on the Dev sets, exploring edit design choices using CAMeLBERT-MSA, AraBERTv02, and ARBERTv2. AraBERTv02 consistently performs best. Subword-level edits, compression, and pruning improve performance, with optimal pruning thresholds of 10 for QALB-2014 and MADAR CODA, and 30 for ZAEBUC. The optimal setup for each dataset (subword, compression, pruning) is presented in Tables 3 and 4. We henceforth refer to this system as sweet (Subword Edit Error Tagger). sweet achieves an F0.5 of 78.8 on QALB-2014 and 86.0 on MADAR CODA, outperforming the Seq2Seq baseline on QALB-2014 and setting a new SOTA on MADAR CODA (though the improvement is not statistically significant). [^3] Statistical significance was done using a two-sided approximate randomization test. On ZAEBUC, it scores 82.7 F0.5, trailing behind the Seq2Seq baseline. Consistent with previous work on text editing Omelianchuk et al. (2020); Straka et al. (2021), we find that iterative correction improves MSA GEC up to two iterations ($\textsc{sweet}^{2}$), achieving 79.3 on QALB-2014 and 83.0 F0.5 on ZAEBUC, with the highest recall on QALB-2014 (70.4). However, iterative correction degrades DA GEC performance. Separating non-punctuation edits ($\textsc{sweet}_{\text{NoPnx}}$) from punctuation edits ($\textsc{sweet}_{\text{Pnx}}$) improves MSA GEC performance. The best setup applies these systems in sequence: two iterations of non-punctuation correction followed by one iteration of punctuation correction ($\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$). This setup achieves the highest F0.5 score among text editing models, setting a new SOTA on QALB-2014 with 80.3. This improvement is statistically significant compared to Seq2Seq++ ($p<0.05$) and is driven by a precision of 83.7. Its performance on ZAEBUC leads other edit tagging techniques but trails behind GPT-4o. For our ensemble models (3-Ensemble), we combine the outputs of the top three non-LLM models per dataset. For QALB-2014 and ZAEBUC, this includes Seq2Seq++, $\textsc{sweet}^{2}$, and the cascaded setup $\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$. For MADAR CODA, we ensemble Seq2Seq++, sweet, and the second-best sweet model using CAMeLBERT-MSA (see Table 10 in Appendix C). The 3-Ensembles outperform single models, achieving SOTA results across all datasets, primarily through increased precision at the cost of recall. Adding GPT-4o’s output to the ensemble further boosts performance (4-Ensemble), reaching an F0.5 of 81.8 on QALB-2014, 86.9 on ZAEBUC, and 88.6 on MADAR CODA. These gains are statistically significant ($p<0.05$) compared to the best baseline and the best non-ensemble model for each dataset.
@@
Tables 5 and 6 present the Test results for MSA and DA GEC, using the best setups identified from the Dev sets. On QALB-2014, the cascaded setup $\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ sets a new SOTA with 80.5 F0.5, outperforming the Seq2Seq and Seq2Seq++ baselines (statistically significant at $p<0.05$). On QALB-2015, this setup matches Seq2Seq++ with an F0.5 of 80.3. Similarly, on ZAEBUC, it achieves 83.1, on par with Seq2Seq++. On MADAR CODA, sweet achieves 86.5, outperforming Seq2Seq++ (though not statistically significant). Our ensemble models further enhance performance across all datasets, reaching 81.7 on QALB-2014, 82.9 on QALB-2015, 87.2 on ZAEBUC, and 88.9 on MADAR CODA. Notably, adding GPT-4o’s output to the ensemble (i.e., the 4-Ensemble) yields statistically significant improvements for MSA GEC. On MADAR CODA, the 3-Ensemble already achieves statistically significant gains, while the addition of GPT-4o does not lead to further improvement.
5.1 Error Analysis
Table 7 presents specific error type performance over the Dev sets of QALB-2014, ZAEBUC, and MADAR CODA. We conduct automatic error analysis using ARETA Belkebir and Habash (2021), an automatic MSA error type annotation tool. ARETA defines error types based on seven classes covering: orthography (O), morphology (M), syntax (X), semantics (S), punctuation (P), merges (Merge-Beginning/Merge-Inside), and splits (Split). On QALB-2014, the cascaded setup $\textsc{sweet}^{2}_{\text{NoPnx}}$ + $\textsc{sweet}_{\text{Pnx}}$ outperforms the Seq2Seq++ baseline on most error types, achieving a macro F0.5 of 71.6. The 4-Ensemble provides a modest improvement, reaching 71.9. On ZAEBUC, while the cascaded sweet setup does not surpass Seq2Seq++ in overall performance (Table 3), it achieves higher scores on most individual error types, with a macro F0.5 of 71.2. This stems from the skewed distribution of error types in the ZAEBUC Dev set, which is dominated by correct words (C) and frequent errors like O and Merge, categories where both models perform similarly. The error types where the cascaded sweet model excels are relatively infrequent (see Table 16 in Appendix G). Notably, while the 4-Ensemble yields the best overall GEC results, it falls short of Seq2Seq++ and the cascaded sweet model in error-type performance, likely due to prioritizing precision over recall (Table 3). On MADAR CODA, neither sweet nor the 4-Ensemble surpasses Seq2Seq++ in error-type performance. While sweet performs slightly better in terms of DA GEC, the improvement is not statistically significant and is driven primarily by precision; in contrast, Seq2Seq++ claims the highest recall (Table 4). The 4-Ensemble further increases precision but at the cost of recall. The high proportion of the unknown (UNK) errors also highlights ARETA’s limitations in capturing dialect-specific errors, as it was primarily designed for MSA.